# Undergraduate Course Allocation

### v1

@@ 3.2 Course Allocation Mechanisms
Using the student utility model, we simulate the outcomes of four mechanisms across 100 random draws of students’ utilities. We assume that students report their utilities truthfully, as all considered mechanisms are either strategy-proof or strategy-proof in the large.
The Pseudo-Market with Priorities (PMP) mechanism. Students are randomly assigned evenly spaced budgets between $1$ and $1+\beta$, with $\beta=\frac{1}{k-1}=0.25$. Students are placed at one of $R=8$ levels of priority in each course. Odd levels 1, 3, 5, and 7 correspond to first, second, third, and fourth-year students ineligible for a course’s reserves, respectively. Even levels correspond to students eligible for reserved seats (see Example 2). Some courses have no students at some levels of priority. For example, if a course has no seat reservations, there are effectively only four levels, one for each year of study.
@@
Our algorithm that searches for a Pseudo-Market Equilibrium with Priorities departs from the one used in Budish (2011) to accommodate priority-specific prices and obtain a tighter bound on the number of over-enrolled courses. As in the proof of Theorem 1, we parameterize prices with $t\in[0,R\overline{b}]^{M}$, where $\overline{b}=1.251>1+\beta$. For each $t\in[0,R\overline{b}]^{M}$, course $c\in{\cal C}$, and priority $r\in{\{1,...,R\}}$, we define priority-specific prices as $p_{c,r}(t)=\max(t_{c}-(r-1)\overline{b},0)$. Then, we look for an equilibrium in this lower-dimensional space.
The algorithm has two phases. Phase I starts with an educated guess for the equilibrium prices based on students’ utility-maximizing course schedules. Then, the algorithm finds each student’s utility-maximizing schedule subject to course prices and adjusts prices proportional to the number of over and under-enrolled seats. The algorithm searches in the lower-dimensional space of $t$-parameters for student demand with a market-clearing error smaller than the theoretical bound $\alpha=\sqrt{kM/2}\approx 43.5$. If the algorithm fails to improve in a given iteration, the current price vector is retained and the size of the price adjustment is reduced for the next iteration. We move to Phase II if the market-clearing error fails to improve within a certain number of iterations, or if the improvement is smaller than $1\%$ when the theoretical bound has been reached. The main innovation of the search process is that course demand is only updated for the subset of students whose demand could be potentially affected by price change on each step. This innovation has proven particularly effective for large student populations, significantly reducing the time required for the search process. [^22] Starting with an educated guess also improves the speed of convergence. In addition, our price adjustment process is a version of the Walras’ tâtonnement procedure, a simpler algorithm than the one used in Othman, Sandholm and Budish (2010). At the same time, we do not consider course complementarities, which is an important part of their paper.
Phase II takes the course prices obtained in Phase I and works to reduce the excess course demand. The phase ensures that no courses are over-enrolled by more than $k-1=4$ seats and that the distribution of over-enrolled courses is smaller than the one observed in the data (see Table 2). [^23] Doing so guarantees that the allocation both has an aggregate market-clearing error below the bound in Theorem 1 and meets the good-by-good bound of Nguyen and Vohra (2022). As a result of Phase II, an allocation with a market-clearing error greater than the theoretical bound $\alpha$ might be obtained. In this case, the algorithm repeats Phases I and II. If there are more than $6$ iterations of Phases I and II, the algorithm restarts with a new set of initial prices. Appendix E.2 provides more details about Phases I and II.
@@
The Deferred Acceptance mechanism with single and multiple tie-breakings. These two mechanisms use the student-proposing deferred acceptance algorithm to allocate course seats to students. The only difference between the two variants is how they break ties between students at the same level of priority in a course. The first variant uses a single tie-breaking order across all courses. To break ties, we use the budgets assigned in the PMP mechanism, placing students with the largest budgets first. We refer to this version of the mechanism as DA-STB. The second variant uses multiple, course-specific tie-breakings. We refer to this version of the mechanism as DA-MTB. Both versions have received a thorough analysis in the many-to-one matching literature (see, e.g., Erdil and Ergin, 2008, 2017). [^24] As students can take more than one course, neither DA-STB nor DA-MTB is strategy-proof (see Roth, 1982). However, Azevedo and Budish (2019) establish that both mechanisms are strategy-proof in the large.
@@
The Random Serial Dictatorship (RSD) with optimal course reserves. Students are ordered by seniority. To break ties, we use the same order as in PMP and DA-STB. Using this order, the mechanism assigns each student the $k$ courses available that offer her the highest utility (or less than $k$, if there are not $k$ such courses). Course seat reservations are treated as described in Section 3.1: for each course and each year of study (including the “All” option), there is a set of departments and a number of reserved seats, which could be zero. If a student is eligible for one of a course’s seat reservations, she can be enrolled in a reserved or a regular seat. Otherwise, the student can only be enrolled in a regular seat. All reserved seats must be occupied before a reserve-eligible student occupies a regular seat.
Using the number of reserved seats observed in the data leads to significant under-performance of RSD with course reserves (see Appendix D.2). The RSD mechanism treats this number of reserves as fixed, resulting in numerous inefficiencies. In practice, the process is more flexible. University departments relax course reservation constraints during the course allocation process. Doing so releases pressure on the department heads to set course reservations precisely at the beginning of the process. In turn, the PMP, DA-STB, and DA-MTB mechanisms compute an effective number of reserved seats in the course allocation process. Hence, comparing the performance of the other three mechanisms to the performance of RSD with a fixed number of reserved seats would not be a fair exercise.
Instead, we estimate the optimal reserved seats for each course in the RSD mechanism. For this purpose, we generate $100$ environments with random student utilities according to the student utility model and a random tie-breaking and run the DA-STB mechanism. To determine the optimal reserved seats for a given course and set of departments, we take the average across all environments (rounded to the nearest integer) of the number of reserve-eligible students assigned to the course. We call the result of the RSD mechanism operated with these calculated reserved seats the RSD with optimal course reserves.
@@ 3.3 Simulation results
Table 5: Each student’s preferred mechanism by year of study.
Pseudo-Market with Priorities Deferred Acceptance with single tie-breaking Deferred Acceptance with multiple tie-breakings
[…]
Year 1 12.9 (1.9) 13.7 (1.7) 3.8 (1.0) 6.7 (1.2) 16.1 (1.2) 18.8 (1.4)
Year 2 24.1 (2.1) 16.7 (1.4) 8.8 (1.2) 7.5 (1.0) 26.4 (1.2) 25.5 (1.1)
Year 3 21.3 (1.8) 12.4 (1.2) 8.5 (1.0) 3.8 (0.7) 22.5 (1.1) 19.9 (1.2)
Year 4 12.9 (1.2) 7.9 (0.9) 3.9 (0.5) 1.6 (0.5) 16.4 (0.9) 14.8 (0.9)
Notes: This table reports the percentage of students who strictly prefer a given mechanism to the benchmark of the Random Serial Dictatorship with optimal course reserves and the percentage of students who strictly prefer the benchmark. Results are averages across $100$ runs with different random component draws.
@@
We evaluate the performance of the Pseudo-Market with Priorities mechanism, Deferred Acceptance with single and multiple tie-breakings mechanisms, and Random Serial Dictatorship mechanism with optimal course reserves across 100 simulation runs. [^25] The program was written in Mathematica. The majority of the simulations were run on a server with 12 Cores of CPU (2.6 GHz of an Intel Xeon Gold 6126 CPU), 128 GBs of RAM, 70 GBs of SSD hard drive space, and Windows Server 2019 Standard operating system. On average, it takes about forty-five minutes to calculate the allocation outcomes of all four mechanisms at one core. Each run corresponds to a random draw of students’ utilities and a random tie-breaking order, which is kept the same across PMP, RSD, and DA-STB. DA-MTB has multiple course-specific tie-breakings.
Student satisfaction. We consider the Random Serial Dictatorship mechanism with optimal course reserves as a benchmark. Table 5 reports the percentage of students who strictly prefer a given mechanism to the benchmark and the percentage of students who strictly prefer the benchmark. The remaining share of students is attributed to indifferences, which account for $48.1\%-94.5\%$ of students across mechanisms and years of study. The large share of indifferences can be explained by students taking under-enrolled courses, as only $11.2\%$ of courses are at or above full capacity in the data.
For second-, third-, and fourth-year students, the share of students who strictly prefer the PMP mechanism to the RSD benchmark is larger than the share of students with the opposite preference. Among students with changing schedules, the mean improvement in utility is $5.1\%$, $8.0\%$, and $9.4\%$, respectively. While slightly more first-year students prefer the benchmark, the PMP mechanism provides an $0.8\%$ mean improvement in utility. This weaker result can be attributed to the gains for more senior students. On average, the PMP mechanism assigns $135$ more course seats to second- through fourth-year students than the RSD benchmark. The DA-STB and DA-MTB mechanisms also perform well relative to the benchmark for second- through fourth-year students. For each DA mechanism, the share of students who strictly prefer it to the RSD benchmark is larger than the share of students with the opposite preference. The difference in these shares is smaller than that for the PMP mechanism for students in each year of study. In general, the efficiency gains for the PMP mechanism are likely understated, as we do not consider any complementarities or more complex student preferences, which the PMP mechanism can easily accommodate.
Table 6: Over-enrolled courses in the Pseudo-Market with Priorities mechanism.
# seats above max capacity $\geq 1$ $\geq 2$ $\geq 3$ $\geq 4$ $\geq 5$
% of courses $2.5$ $0.4$ $0.1$ $0.02$ $0.0$
In principle, the PMP mechanism may be delivering better outcomes to students due to the market-clearing error, which allows some courses to be over-enrolled. [^26] Alternatively, we could have compared the performance of the four mechanisms with quotas equal to the number of seats assigned in the PMP mechanism for each run. We thank Olivier Tercieux for this suggestion. Table 6 reports the distribution of over-enrollment in courses. In line with the good-by-good clearing bound of Nguyen and Vohra (2022), no courses are over-enrolled by more than $k-1=4$ seats. In addition, the percent of over-enrolled courses is smaller than that observed in the university’s final allocation. Less than $2.5\%$ of courses are over-enrolled by more than one seat, as opposed to $7.3\%$ in the data. The average market-clearing error across outcomes is about $21.4$ seats, less than half of the worst-case bound $\alpha\approx 43.5$.
Table 7: The change in the standard deviation of students’ utilities by year of study.
Pseudo-Market with Priorities Deferred Acceptance with single tie-breaking Deferred Acceptance with multiple tie-breakings
Year 1 -3.3% -0.6% -2.4%
Year 2 -5.6% -1.5% -4.0%
Year 3 -3.4% -0.2% -1.9%
Year 4 -1.1% 0.2% -0.7%
@@
Notes: This table reports the percent change in the standard deviation of students’ utilities compared to the benchmark of to the Random Serial Dictatorship with optimal course reserves for three mechanisms. Results are averages across $100$ runs with different random component draws.
@@
Allocation fairness. We compare allocation fairness across mechanisms using two metrics. Table 7 shows that, within each year of study, the PMP and DA-MTB mechanisms decrease the standard deviation of students’ utilities relative to the RSD benchmark. The same result holds for first- through third-year student in the DA-STB mechanism. For each year of study, this finding is the strongest for the PMP mechanism. Between the two DA mechanisms, the DA-MTB mechanism performs slightly better. For each mechanism, the largest reduction occurs for second-year students. This is in line with Table 5, where Year 2 had the fewest percent of students who are indifferent between each mechanism and the benchmark.
@@
Table 8 reports results on students who experience schedule envy towards students of weakly lower priority in each course. The first column shows the percentage of students with no such envy. The values in each subsequent column are obtained by removing one course from the envied schedules and checking whether the student still experiences envy. In line with the result of Theorem 4, in the PMP mechanism, no students experience schedule envy greater than a single course toward students of weakly lower priority in each course. None of the other three mechanisms satisfy this property. About $9.1\%$ of students experience envy in the PMP mechanism, substantially less than in the RSD, DA-STB, and DA-MTB mechanisms ($17.2\%$, $14.1\%$, and $13.3\%$, respectively). [^27] Ours is the first paper to compare DA-STB and DA-MTB in terms of the envy students experience to others of a weakly lower priority. See Abdulkadiroğlu, Pathak and Roth (2009) and Ashlagi, Nikzad and Romm (2019) for a comparison of these mechanisms using other measures of fairness and efficiency. Relative to the RSD benchmark, the PMP mechanism reduces the percentage of students who experience any envy toward students of weakly lower priority by 8 points, or almost $500$ students.
@@
Notes: This table reports the percentage of students who prefer the schedule of a student with weakly lower priority in each course in each of the four mechanisms. Results are averages across 100 runs with different random component draws. The first column includes students who experience no envy. The other columns display students who experience schedule envy bounded by $1,...,5$ courses.
@@
We can also study allocation fairness between priority levels. In the PMP, DA-STB, and DA-MTB mechanisms, no student can ever improve her course schedule by taking a seat that is occupied by a student of strictly lower priority. However, in the RSD mechanism, $16.2\%$ of students can. These instances occur when a course reserves too few of its seats for a designated group of students. Then, a reserve-eligible student late in the tie-breaking order may miss out on a seat when a reserve-ineligible student obtains one.
@@
Limitations. Our student utility model has limitations that a richer data set and a more detailed empirical analysis could address. Most notably, the additive utility specification may be too restrictive, as it fails to consider some complementarities and substitutabilities across courses. For example, courses covering similar content may be substitutes, while a course requiring another as a corequisite (i.e., to be taken at the same time) may be complements. Still, the additive structure of student preferences allows us to compare the performance of our new mechanism with the DA-STB and DA-MTB mechanisms. Unlike these mechanisms, the PMP mechanism can easily be extended to more general preference specifications.
@@
A first step in expanding this analysis beyond additive utilities is to consider preferences over course sections rather than courses. Universities often offer the same course in different time slots or with different instructors. A section-level analysis would relax the assumption of additive utility, as each student cannot be enrolled in two sections of the same course. In the data, many courses have several sections, with some being much more popular than others. Since we aggregate enrollment and capacities across sections, if one of a course’s sections is not at maximum capacity, the course cannot be either. Overall, $30\%$ of course sections are at maximum capacity, in comparison to only $11.2\%$ of courses. Consequently, we expect the gains from the PMP mechanism presented earlier in this section to be understated.
@@
Another approach would be to maintain the additive utility specification, but consider more factors that influence student preferences. Our student utility model assumes that a student’s preferences depend only on her college and year of study. Allowing utilities to depend on a student’s department would capture additional heterogeneity in preferences. Additionally, information about past enrollment would help to disentangle preferences over courses from constraints based on the course’s prerequisites. This information would also allow us to pin down students’ choice sets. However, past enrollment, along with other characteristics (e.g., demographics) are not available in the data.
We conclude this section by summarizing the main insights from our empirical analysis. The Pseudo-Market with Priorities mechanism delivers a substantially fairer outcome than the Random Serial Dictatorship with optimal course reserves, both in terms of the standard deviation of student utilities and the percentage of students with envy toward other students of weakly lower priority. Furthermore, student satisfaction improves for second- through fourth-year students. The two variants of the DA mechanism share some benefits of the PMP mechanism, but have weaker fairness results and cannot be extended to accommodate more general student preferences. Appendix D presents additional simulation results, detailing the equilibrium prices in the PMP mechanism (Section D.1), the performance of six alternative mechanisms (Section D.2), results under an alternative priority structure (Section D.3), and robustness checks with respect to the size of student choice sets and the standard deviation of the noise parameter (Section D.4).

### v2

@@ 3.2 Course Allocation Mechanisms
We analyze the four mechanisms described below. We assume students report their utilities truthfully, as each mechanism is strategy-proof or strategy-proof in the large.
@@
The Pseudo-Market with Priorities (PMP) mechanism. Students are randomly assigned evenly spaced budgets between $1$ and $1+\beta$ (with $\beta=\frac{1}{k-1}=0.25$) and placed at one of $R=8$ levels of priority in each course. As in the proof of Theorem 1, we parameterize prices with $t\in[0,R\overline{b}]^{M}$, where $\overline{b}=1.251>1+\beta$. For each $t\in[0,R\overline{b}]^{M}$, course $c\in{\cal C}$, and priority level $r\in{\{1,...,R\}}$, we define priority-specific prices as $p_{c,r}(t)=\max(t_{c}-(r-1)\overline{b},0)$. Then, we look for an equilibrium in the lower-dimensional $t$-space.
The search process has two phases. Phase I starts with an educated guess for course prices based on students’ utility-maximizing course schedules. The algorithm searches in the lower-dimensional space of $t$ parameters for a market-clearing error smaller than the theoretical bound $\alpha=\sqrt{kM/2}\approx 43.5$, iteratively computing each student’s utility-maximizing schedule and adjusting prices proportional to the number of over and under-enrolled seats. We move to Phase II if the market-clearing error fails to improve within a certain number of iterations or if the improvement is smaller than $1\%$ after the theoretical bound has been reached. Phase II takes the Phase I prices and works to reduce excess demand, ensuring that no course is over-enrolled by more than k1=4k-1=4 seats and the distribution of over-enrolled courses is smaller than the one observed in the data (see Table 2). [^20] Doing so guarantees the allocation has a market-clearing error below the bound in Theorem 1 and meets the good-by-good bound of Lin et al. (2022). A proof on the existence of an equilibirum with the good-by-good bound in our setting is available by request. We return back to Phase I if an allocation with a market-clearing error greater than $\alpha$ is obtained in Phase II.
@@
The Deferred Acceptance mechanism with single and multiple tie-breakings. These mechanisms use the student-proposing deferred acceptance algorithm to allocate course seats and differ in how they break ties between students at the same level of priority in a course. The first variant uses a single tie-breaking order across all courses. To break ties, we use the budgets assigned in the PMP mechanism, placing students with larger budgets earlier. Without any course seat reservations, this mechanism is equivalent to a random serial dictatorship (with students ordered by seniority and ties broken randomly) and can lead to significant envy between students. The second variant uses multiple, course-specific tie-breakings. We refer to these mechanisms as DA-STB and DA-MTB, respectively. Both have received a thorough analysis in the many-to-one matching literature, but their many-to-many counterparts are underexplored (see, e.g., Erdil and Ergin, 2008, 2017).
@@
The Random Serial Dictatorship (RSD) with optimal course reserves. This mechanism orders students by seniority, breaking ties using the same order as in PMP and DA-STB, and assigns each student the (up to) $k$ courses available that offer her the highest utility. Course seat reservations are treated as described in Section 3.1. If a student is eligible for one of a course’s seat reservations, she can enroll in a reserved or regular seat. Otherwise, the student can only be enrolled in a regular seat. Reserved seats must be occupied before a reserve-eligible student occupies a regular seat.
Using the number of reserved seats observed in the data leads to significant under-performance of RSD with course reserves (see Appendix D.2). In practice, the university’s registration system has two phases. The first is a centralized admission process, which uses a version of the random serial dictatorship with course reserves. The second is a decentralized stage, where departments relax reserves to fill unoccupied seats with waitlisted students. This second phase is not well-structured. Importantly, the university administration sets course reserves more generously in the first phase than it would if the second phase did not exist. This guarantees that prioritized students have access to course seats but leaves many unfilled reserved seats to be assigned in the second phase. Hence, comparing the performance of the other three mechanisms to the performance of RSD with these reserved seats would not be a fair exercise. Instead, we estimate the optimal reserved seats by generating $100$ environments with random student choice sets and utilities, setting a random tie-breaking order, and running the DA-STB mechanism. We take the average number of reserve-eligible students assigned to each course across all environments (rounded to the nearest integer) as the optimal number of reserved seats. We call the result of the RSD mechanism operated with these estimated reserved seats the RSD with optimal course reserves. [^21] The term “optimal” is not used here in the sense of maximizing some objective, but rather to denote a reasonable and practical specification of reserve prices within the current setting.
In the next section, we present simulation results comparing the PMP, DA-STB, DA-MTB, and RSD mechanisms across five measures. These include two measures of student satisfaction: (i) each student’s preferred mechanism and (ii) mean student utility; and three measures of allocation fairness: (iii) the standard deviation of students’ utilities, (iv) the amount of envy (toward students with weakly lower priority), and (v) priority violations (students desiring a course being taken by a student of strictly lower priority). From theory alone, we know that no student can experience envy greater than a single course in the PMP mechanism (Theorem 4), and that the PMP, DA-STB, and DA-MTB mechanisms all prevent priority violations. Otherwise, it is not clear how the four mechanisms will perform relative to one another. Our simulations provide comparisons for the remaining measures and allow us to see if the envy and priority violations have important magnitudes.
@@ 3.3 Simulation results
We evaluate the performance of the Pseudo-Market with Priorities, Random Serial Dictatorship with optimal course reserves, and Deferred Acceptance with single and multiple tie-breakings mechanisms across 100 simulation runs. [^22] Simulations were run on Mathematica with a server with 12 Cores of CPU (2.6 GHz Intel Xeon Gold 6126 CPU), 128 GBs of RAM, 70 GBs of SSD hard drive, and a Windows Server 2019 Standard operating system. It takes about forty-five minutes to calculate the outcomes of all four mechanisms for each run. Each run corresponds to a random draw of student choice sets, utilities, a single tie-breaking order kept the same across PMP, DA-STB, and RSD, and multiple course-specific tie-breakings for DA-MTB.
Figure 1: Each student’s preferred mechanism by year of study.
@@
Notes: This figure reports the percent of students who strictly prefer each given mechanism to the Random Serial Dictatorship with optimal course reserves benchmark and the percent of students who strictly prefer the benchmark. Results are averages across $100$ runs with different random component draws.
Student satisfaction. We take the RSD with optimal course reserves as a benchmark. Figure 1 displays results on the percent of students who strictly prefer each given mechanism to RSD and the percent of students who strictly prefer RSD. Bars represent shares of the total student population and are broken down by year of study. The large share of indifferences across mechanisms can be explained by students taking under-enrolled courses, as only $11.2\%$ of courses are at or above capacity in the data. Around $17.9\%$ of students prefer PMP to RSD, as opposed to $12.8\%$ of students preferring RSD to PMP. As a result, over $5\%$ more of the population - or $300$ more students - strictly prefer the PMP mechanism. A greater number of first-year students strictly prefer the benchmark to each alternative mechanism. The weaker result for Year 1 students can be attributed to the gains for more senior students, who, on average, are assigned $135$ more course seats under PMP than RSD.
In contrast, $88.7\%$ of students are indifferent between the DA-STB and RSD mechanisms. The vast majority of students being indifferent between the two mechanisms can be explained by our using DA-STB to select the optimal course reserves for the RSD mechanism. Moreover, the outcomes of RSD and DA-STB coincide in the absence of course reserves, and if both mechanisms employ the same tie-breaking rule. DA-MTB results in only $59.7\%$ of students with indifferences, with $20.4\%$ of students strictly preferring DA-MTB and $19.9\%$ strictly preferring RSD. These results are a direct consequence of DA-MTB breaking ties on a per-course basis, producing both winners and losers among students. Also, direct calculations show that more students strictly prefer the PMP mechanism to each DA mechanism in each year of study. [^23] The results of the direct comparison of DA-STB and DA-MTB with PMP are available upon request.
Figure 2: Improvements in utility among students with changing schedules.
Notes: This figure reports the percent improvement in mean utility among students with changing schedules for each given mechanism over the Random Serial Dictatorship with optimal course reserves benchmark. On average, the PMP mechanism produces a different schedule for 416, 658, 478, and 298 students in Years 1, 2, 3, and 4, respectively; the DA-STB mechanism does so for 164, 263, 174, and 78 students; and the DA-MTB mechanism does so for 547, 836, 603, and 444 students. Results are averages across $100$ runs with different random component draws.
Figure 2 presents changes in cardinal utility, displaying each mechanism’s percent improvement in mean utility over the RSD mechanism among students with changing schedules. The PMP mechanism improves mean utility for all years of study, with larger gains for students in later years of study. In both DA mechanisms, the percent improvement in mean utility is negative for Year 1 students and close to zero for Year 2 students. Among students with changing schedules, the DA-STB mechanism provides the largest improvement in mean utility for Years 3 and 4. However, the DA-STB mechanism changes the schedules of only 174 and 78 students for Years 3 and 4, respectively. In contrast, the PMP mechanism changes the schedules for 479 and 298 students for Years 3 and 4. DA-MTB delivers almost no improvement in mean utilities for each cohort. For each year of study, across all students, the PMP mechanism has the highest mean utility of the four mechanisms.
@@
Allocation fairness. We compare allocation fairness across mechanisms using three metrics. First, Figure 3 shows that, within each year of study, the PMP and DA-MTB mechanisms decrease the standard deviation of students’ utilities relative to the RSD benchmark. The same result holds for all but fourth-year students in the DA-STB mechanism. For each year of study, this finding is the strongest for the PMP mechanism. For each mechanism, the largest reduction occurs for second-year students.
Figure 3: The change in the standard deviation of students’ utilities.
Notes: This figure reports the percent change in the standard deviation of students’ utilities for each given mechanism compared to the Random Serial Dictatorship with optimal course reserves benchmark. Results are averages across $100$ runs with different random component draws.
@@
Second, Table 5 reports results on schedule envy toward students of weakly lower priority in each course. The first column displays the percent of students without envy. The values in each subsequent column are obtained by removing one course from each envied schedules and checking whether the student still experiences envy. In line with our theoretical results for the PMP mechanism (see Theorem 4), no student has envy greater than a single course toward a student of weakly lower priority. As Table 5 shows, this property is not satisfied by the other three mechanisms. About $9.1\%$ of students experience envy in the PMP mechanism, substantially less than in the RSD, DA-STB, and DA-MTB mechanisms ($17.2\%$, $14.1\%$, and $13.3\%$, respectively). Relative to the RSD benchmark, $8.1\%$ less of the total population (or almost 500 students) experience any envy in the PMP mechanism.
@@
Notes: This table reports the percentage of students who prefer the schedule of a student with weakly lower priority in each course in each of the four mechanisms. The first column includes students who experience no envy. The other columns display students who experience schedule envy bounded by 1,,51,...,5 courses. Results are averages across 100 runs with different random component draws.
@@
Third, we examine fairness between levels of priority. The PMP mechanism is approximately stable, guaranteeing that no student can improve her schedule by taking a seat assigned to a student with strictly lower priority in the course (see Theorem 2). As students’ preferences are additive over courses, the DA-STB and DA-MTB mechanisms also ensure that no priorities are violated. However, in the RSD mechanism, $16.2\%$ of students would benefit from taking a seat assigned to a student with strictly lower priority. These instances occur when a course reserves too few of its seats, meaning a reserve-eligible student selecting courses late in the RSD mechanism loses out on a seat when a reserve-ineligible student selecting earlier obtains one.
Our simulations present important comparisons that would be difficult to derive from theory. While there is some variation across years of study, the PMP mechanism generally outperforms the other three mechanisms in each measure of student satisfaction and allocation fairness. [^24] In principle, it may be delivering better outcomes due to the market-clearing error, which allows some courses to be over-enrolled. The average market-clearing error across outcomes is $\approx 21.4$, less than half of the worst-case bound of $\approx 43.5$. In line with the bound in Nguyen and Vohra (2022), no course is over-enrolled by more than k1=4k-1=4 seats. About $3\%$ of courses are also over-enrolled, as opposed to $7.3\%$ in the data. The comparison of DA-STB and DA-MTB in many-to-many settings is not readily available in the literature. Hopefully, these results will direct further theoretical investigations of these mechanisms. Appendix D contains additional simulation results, including a description of the equilibrium prices in the PMP mechanism (Section D.1), a comparison with six alternative mechanisms (Section D.2), results under an alternative priority structure (Section D.3), and further robustness checks (Section D.4).
@@
Limitations. Our student utility model has limitations that a more detailed empirical analysis could address. Notably, assuming additive utilities may be too restrictive, as it fails to consider some complementarities and substitutabilities across courses. Courses covering similar content are substitutes, while a course requiring another as a corequisite (i.e., to be taken at the same time) are complements. Additive utilities allow us to compare the performance of the PMP mechanism with the DA-STB and DA-MTB mechanisms. Unlike these mechanisms, the PMP mechanism can be extended to more general preferences.
@@
One step in expanding this analysis beyond additive utilities is to consider preferences over course sections rather than courses. Universities offer the same course at different times or with different instructors, restricting students from enrolling in two sections of the same course. In the data, many courses have several sections, with some much more popular than others. Since we aggregate enrollment and capacities across sections, if one of a course’s sections is not at maximum capacity, the course cannot be either. $30\%$ of sections are at maximum capacity, as opposed to only $11.2\%$ of courses. Consequently, we expect the gains from the PMP mechanism presented earlier in this section to be understated.
@@
Alternatively, we could consider additional factors that influence preferences. Our utility model assumes preferences only depend on a student’s college and year of study. Allowing utilities to depend on a student’s department would capture more heterogeneity in preferences. Information about past enrollment would disentangle preferences over courses from constraints based on the course’s prerequisites, pinning down student choice sets. Past enrollment, along with other characteristics (e.g., demographics), are not available in our data.